Find the normal approximation for the binomial probability that x = 5, where n = 12 and p = 0.6. Compare this probability to the value of P(x=5)
Can someone help me

Answer by stanbon(75887)

Find the normal approximation for the binomial probability that x = 5,
where n = 12 and p = 0.6. Compare this probability to the value of P(x=5)
---------------------
u = 0.6*12 = 7.2
s = sqrt(7.2*0.4) = 1.6971
P(x=5) = P(4.5< x <5.5) = normalcdf(4.5,5,5,7.2,1.6971) = 0.1024
---
P(x=5) = binompdf(12,0.6,5) = 0.1009
=============
Note: I used a TI-84 to do the calculating.
================
Cheers,
Stan H.